# Constraints on SMEFT operators from $Z\to\mu\mu bb$ decay

### v1

@@ VI Conclusions
In this work, we have presented a phenomenological study of the $Z\to\mu\mu bb$ final state as a probe of physics beyond the Standard Model within the framework of the Standard Model Effective Field Theory. The analysis is entirely based on Monte Carlo simulations and aims at assessing the sensitivity of this channel to a selected set of dimension-six SMEFT operators.
Signal and relevant background processes were generated at leading order using MadGraph5_aMC@NLO, followed by parton showering and hadronization with Pythia, and a fast detector simulation with Delphes. A realistic event selection strategy was implemented, including trigger requirements, lepton isolation, $b$-jet identification, and kinematic separation cuts. After event reconstruction, the invariant mass distribution of the $\mu\mu bb$ system, together with transverse momentum spectra of the final-state objects, was used to validate the signal modeling and background suppression.
To quantify possible deviations from the Standard Model, we employed the SMEFT reweighting technique using a UFO implementation of dimension-six operators. For each operator under consideration, event weights corresponding to multiple values of the WC were extracted at the generator level and propagated through the full analysis chain. The total event yield after selection was shown to follow a quadratic dependence on the WC, reflecting the interference and pure SMEFT contributions to the cross section.
A likelihood-based statistical analysis was performed by constructing a test statistic based on the expected event yields. Assuming the Standard Model prediction as the null hypothesis, we derived 95% confidence level constraints on the WCs by identifying the parameter values corresponding to the critical value of the test statistic. This procedure provides a transparent and robust estimate of the sensitivity of the $\mu\mu bb$ channel to SMEFT effects, in the absence of experimental systematic uncertainties associated with real data.
The results demonstrate that the $Z\to\mu\mu bb$ final state constitutes a viable and complementary probe of electroweak and Higgs-related effective operators. Although the present study is limited to leading-order simulations and an inclusive cut-based analysis, it establishes a solid baseline for future improvements. Possible extensions include the incorporation of higher-order corrections, a differential shape analysis, the inclusion of systematic uncertainties, and a global fit combining multiple final states.
Overall, this work highlights the potential of precision studies of multi-body $Z$-boson decay final states within the SMEFT framework and demonstrates that mixed leptonic–hadronic channels can provide unique and complementary information on flavor-resolved new physics interactions.

### v3

@@ VI Conclusions
In this work, we have investigated the $\mu^{+}\mu^{-}b\bar{b}$ final state in the reconstructed $Z$-pole region as a probe of selected SMEFT operators. The analysis focuses on flavor-resolved interactions involving second-generation leptons and third-generation quarks, including both four-fermion operators and operators that modify effective ZZ–fermion couplings.
Signal and relevant background processes were simulated at leading order using MadGraph5_aMC@NLO, followed by parton showering and hadronization with Pythia8, and fast detector simulation with Delphes. The event selection includes muon trigger requirements, lepton isolation, $b$-jet identification, missing-energy suppression, and angular separation cuts. The reconstructed invariant mass of the $\mu^{+}\mu^{-}b\bar{b}$ system is used to define a signal region around the $Z$ pole and to extract constraints on possible SMEFT effects.
The dependence of the selected event yield on individual Wilson coefficients was obtained using the SMEFT reweighting procedure. For each operator, the yield was parametrized as a quadratic function of the Wilson coefficient, including both the SM–dimension-six interference and the quadratic dimension-six contribution. Using an Asimov likelihood-ratio method, we derived expected 95% C.L. constraints on the selected operators at integrated luminosities of $138~\mathrm{fb}^{-1}$ and $3000~\mathrm{fb}^{-1}$.
The results show that the $\mu^{+}\mu^{-}b\bar{b}$ final state can provide useful channel-specific constraints on flavor-resolved SMEFT interactions, especially on operators involving muons and bottom quarks. These constraints are complementary to those obtained from inclusive electroweak measurements, purely leptonic $Z$ decays, and global SMEFT analyses. The projected HL-LHC luminosity further improves the reach, illustrating the potential value of including mixed leptonic–hadronic $Z$-pole final states in future SMEFT studies.
The present analysis provides a baseline phenomenological study of this channel. Several refinements can be incorporated in future work, including higher-order QCD and electroweak corrections, scale and PDF uncertainties, experimental systematic uncertainties, correlations among multiple Wilson coefficients, SMEFT-induced modifications of the total $Z$ width, and a more complete treatment of electroweak interference effects in the irreducible $\mu^{+}\mu^{-}b\bar{b}$ final state. These extensions would allow the constraints from this channel to be embedded more directly into global SMEFT fits and precision electroweak interpretations.